OpenStudy (anonymous):

Evaluate the limit. (Use symbolic notation and fractions where needed.) \[\lim\limits_{x \to \infty} [\frac{1}{x}-\frac{6}{x+53}] = \]

hartnn (hartnn):

hi @rjo89 :) try substuting, x = 1/y then 1/x =... ? and since x->infinity, 1/x -> ....?

OpenStudy (anonymous):

thanks! the answer's 1

hartnn (hartnn):

oh, is it! ?? i got the answer as 0 how did u get it as 1 ?

OpenStudy (tkhunny):

\(\dfrac{1}{x} - \dfrac{6}{x+53} = \dfrac{x+53 - 6x}{x(x+53)} = \dfrac{53 - 5x}{x^{2} + 53x}\) And thus we see rather clearly that the limit as x increases without bound is 0. I, too, would be quite interested to see how '1' was obtained.